OpenStudy (secret-ninja):

Medals. ^-^ What is the equation of a line that goes through the points (58.5, 59) and (73, 74)? Show your work.

OpenStudy (secret-ninja):

@surjithayer @nincompoop Do you know this?

OpenStudy (anonymous):

\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\left( x-x _{1} \right)\]